OpenStudy (anonymous):

anyone proove 0/0=20???

OpenStudy (anonymous):

undefined cant 0/0

terenzreignz (terenzreignz):

No.. unless you allow division by zero... in which case, actually, 0/0 is any number.

OpenStudy (unklerhaukus):

\[0/0=20\\\quad0=20\times0\\\quad0=0\\\qquad\quad\boxed{}\]

OpenStudy (anonymous):

as we know formula a^a-b^b=(a+b)(a-b) so((10^10)-(10^10))/(10-10) so.....((10+10)*(10-10))/(10-10)=20